OpenStudy (anonymous):

If there are 10 spa beds and the rate of occupying the spa beds is 0.8. The state will force the spa center to shut down if not more than 2 beds are filled. What would be the probability for the spa center to go out of business?

OpenStudy (tkhunny):

p(not more that 2) = 1 - p(0 or 1 or 2) p(0) = 0.2^10 p(1) = 10*.2^9*.8^1 p(2) = You tell me.

OpenStudy (anonymous):

p2= 10!/(2! *(10-2)!) * 0.8^ 2 * (1-0.8)^8

OpenStudy (tkhunny):

Super. Now, did you answer the question?

OpenStudy (anonymous):

yes.